find the value of x in the equation 9-4x=57
Answer by tutorcecilia(2152)

9-4x=57
9-9-4x=57-9
-4x=48
-4x/-4=48/-4
x=-12
.
Check by plugging x=-12 back into the original equation and solve:
9-4x=57
9-4(-12)=57
57=57 [checks out]